Let f(x)=x^2 and let g(x)=(x-3)^2 +2
a.) Give the formula for g in terms of f, and describe the relationship between f and g in words.
b.) Is g a quadratic function? If so, find its standard form and the parameters a, b, and c.

Answer by stanbon(75887)

Let f(x)=x^2 and let g(x)=(x-3)^2 +2
a.) Give the formula for g in terms of f, and describe the relationship between f and g in words.
g[f(x)] = g[x^2] = (x^2-3)^2 + 2
---------------------------------------
Describtion: Translate f 3 units to the right and raise each point 2 units.
-----------------------------------------------------------
b.) Is g a quadratic function? If so, find its standard form and the parameters a, b, and c.
Ans: Yes, g is quadratic.
g(x) = x^2-6x+9+2 = x^2-6x+11
a = 1
b = 6
c = 11
------------
Cheers,
Stan H.